a) Give parametric equation (point-direction form) of the line which lies on both of the planes:

x + y + z = 1 and −x + 2y + 10z = 2. What is the direction d of this line?

b) Let n1 and n2 be the normal vectors to the two given planes. Without actual computation,

describe the relationship between d and n1 × n2.


1
Expert's answer
2021-09-28T14:38:13-0400

Let "y=0" , then:

"11z=3\implies z=3/11"

"x=8\/11"


The direction vector of this line:

"\overline{d}=\overline{n}_1\times \overline{n}_2=\begin{vmatrix}
 i & j&k \\
 1 & 1&1\\
-1&2&10
\end{vmatrix}=8i-11j+3k"


The equation of the line:

"\frac{x-8/11}{8}=\frac{y}{-11}=\frac{z-3/11}{3}"


Parametric equation of the line:

"\begin{cases}
 x=p_1t+x_0 \\
 y=p_2t+y_0\\
 z=p_3t+z_0
\end{cases}"


"\begin{cases}
 x=8t+8/11 \\
 y=-11t\\
 z=3t+3/11
\end{cases}"


"|\overline{d}|=|\overline{n}_1\times \overline{n}_2|=\sqrt{8^2+11^2+3^2}=\sqrt{194}"

direction of the line:

"cos\alpha=8/\sqrt{194}"

"cos\beta=-11/\sqrt{194}"

"cos\gamma=3/\sqrt{194}"